# Trajectory Planning for Autonomous Vehicle Using Iterative Reward Prediction in Reinforcement Learning

### v3

@@ III-D Uncertainty Propagation
During RP and IRP, the expected future states are considered; however, the variance and uncertainty of the predicted future states are not considered. Note that not considering the variance of the prediction leads to inaccurate function approximation and the RL agent being unaware of the uncertainty. In addition, ignoring the uncertainty can cause the agent to perform sudden deceleration or become involved in an accident. For example, if a low-functioning controller is employed to track a trajectory, and the RL agent does not take that into account, then the AV may encounter various dangerous situations and the predicted rewards may have significant error. To consider the variance of the prediction and the uncertainty of future states, we utilize uncertainty propagation on the RP and IRP motivated by [21, 22].
The uncertainty propagation process is built on the Kalman filter; however, the measurement update process of the Kalman filter is removed because observing future states is impossible. The uncertainty propagation utilized on the RP is expressed as follows:
@@
where $\Sigma^{e}_{t,\tau}$, and $\Sigma^{o}_{t,\tau}$ represent the covariance of the Gaussian random variables $\check{s}^{e}_{t,\tau}$, and $\check{s}^{o}_{t,\tau}$, respectively, $F$ is the state transition matrix and $Q^{e}_{\tau}$, $Q^{o}_{\tau}$ represent the process noise of the AV and the other traffic participants, respectively, where $Q^{e}_{\tau}$ is attributed to localization, control error, and $Q^{o}_{\tau}$ is attributed to object detection, trajectory prediction error respectively. In addition, $\check{s}^{e}_{t,\tau}$ and $\check{s}^{o}_{t,\tau}$ are used in IRP to perform collision checking. Here, the ellipse defined by the covariance matrix can provide an upper bound of the probability $1-\delta$ that the AV and other traffic participants exist. However, to check for a collision between the AV and other traffic participants, the rectangle shapes of the AV and other traffic participants must be considered. In the proposed method, We compute the Minkowski sum of the rectangle and the ellipse [21]. The new shape from the Minkowski sum is then utilized for collision checking, which guarantees probability of $(1-\delta)^{2}$ whether collide or not. The uncertainty propagation and collision checking process using the Minkowski sum are illustrated in Fig.5
Figure 5: Predicted states of the AV(red) and another traffic participant(blue). The state of the AV and the goal are represented as a red circles. The predicted states of the other traffic participant and the AV are represented as blue boxes and red boxes for, respectively. Their covariance is represented as the dark blue and dark red circles in the center of the corresponding boxes, respectively. Note that a larger covariance results in a larger circle. The Minkowski sum of each state and the covariance are shown in light blue and light red rounded rectangle, respectively.
@@ III-E Overall Algorithm
The proposed method is based on existing deterministic policy gradient algorithms [26, 20]. We modified the critic update process by utilizing the IRP, and the uncertainty propagation method. The pseudocode of the proposed method is given in Algorithm 1. Here, in lines 6-7, the uncertainty propagation and Minkowski sum are executed using the state of the AV and the other traffic participants $s^{e}_{t+\tau,0},s^{o}_{t+\tau,0}$. In line 9, the state $s_{t+\tau,\delta}$ is predicted by the agent’s predicted action $\pi(s_{t+\tau,0})$, where $s_{t+\tau,0}$ comprises the predicted AV’s state $s^{e}_{t+\tau-\delta,\delta}$ and the other traffic participants’ predicted state $s^{o}_{t,\tau+\delta}$ from the trajectory prediction module at time $t$. In line 10, the reward is predicted during the transition from $s_{t+\tau,0}$ to $s_{t+\tau,\delta}$.
@@
3:     $r^{\prime}\leftarrow r_{t,\delta}$triangleright Initialize predicted reward with received reward from the simulator
@@
10:         $r\leftarrow g(s,s^{\prime})$triangleright Prediction of $r_{t+\tau,\delta}$
11:         $r^{\prime}\leftarrow r^{\prime}+\gamma^{\tau/\delta}r$triangleright Update Predicted Reward
12:         $s^{e},s^{o}\leftarrow f^{-1}(s^{\prime})$triangleright Update Next State
@@
17:     $s^{\prime}\leftarrow f({s^{e}}^{\prime},{s^{o}}^{\prime})$triangleright Merge States
18:     Set $y_{t}=r^{\prime}+\gamma^{T/\delta}Q(s^{\prime},\pi(s^{\prime}|\theta^{\pi})|\theta^{Q})$
@@ IV Experiments
The proposed method and baseline methods were evaluated using the CARLA simulator. The experimental configuration, baseline methods, and implementation details are described in the following.
@@ IV-A Experiment Configuration
We evaluated the proposed and baseline methods in four distinct scenarios. Scenario 1 involved a lane following scenario with static obstacles, scenario 2 involved a lane following scenario with traffic participants, scenario 3 involved a lane changing scenario with traffic participants, and scenario 4 is an overtaking parked cars scenario with traffic participants. All necessary inputs, e.g., route path, object detection, trajectory prediction, and localization were given. Success was determined if the AV reached the goal without a collision within specified time. In this evaluation, the goal was $130m$ ahead of the AV’s initial position. Here, a maximum lateral deviation of $1.5m$ from the center of the target lane was permitted.
In each scenario, the AV was spawned on the road with a random lateral deviation of $[-1.5m,1.5m]$ from the center of the road, and a random heading angle deviation of $[-20deg,20deg]$ with random initial speed of $[5km/h,15km/h]$. In scenario 1, a maximum of two static obstacles(i.e., vehicles) were spawned at random positions with a lateral deviation of $[-0.5m,0.5m]$ from the center of the road, and a heading angle deviation of $[-20deg,20deg]$. Scenario 2 included a maximum of five randomly spawned traffic participants. In scenario 3, the other traffic participants were the same as in scenario 2; however, the AV’s goal was to change lanes. In scenario 4, a maximum of two static obstacles(parked cars) and three traffic participants were spawned, and the goal of the AV was to overtake the parked cars while avoiding collisions with the other traffic participants. Note that the traffic participants in each scenario were designed to change lanes randomly.
@@ IV-B Baseline Methods
The following two baseline methods were considered in this evaluation. In baseline 1, the output of an agent was a control command, and the input feature was the same as the proposed method. In baseline 2, the output of the agent was the goal of a trajectory which was the same as the proposed method without RP, IRP, and uncertainty propagation. In addition, we evaluated our methods individually as follows: RP, IRP, and IRP with uncertainty propagation.
@@ IV-C Implementation Details
All five methods, i.e., baseline1, baseline2, RP, IRP, and IRP with uncertainty propagation were implemented using the DDPG algorithm.
@@ IV-C1 State
The features of the state space $s$ are composed of $s^{e}$ and $s^{o}$. Here, $s^{e}$ is the state of the AV and comprises $(d,\dot{d},\ddot{d},\dot{\sigma},\ddot{\sigma},\theta,v_{speed\_limit})$, where $(\sigma,d)$ is the longitudinal and lateral position on the Frenet frame, $\theta$ is the heading angle difference with the center of the road, and $v_{speed\_limit}$ is the speed limit of the road. $s^{o}$ is composed of $s^{o,k},k\in\mathbb{N}$, where $\mathbb{N}$ is a natural number. $s^{o,k}$ is the state of the $k$th traffic participant, which is composed of $(\sigma,d,\theta{{}^{\prime}},l,w,v_{\sigma},v_{d})$, where $(\sigma,d)$ is the position, $\theta{{}^{\prime}}$ is the heading angle difference, $(v_{\sigma},v_{d})$ is the velocity of the $k$th traffic participant on the Frenet frame, and $(l,w)$ represents the length and width of the vehicle in consideration of the Minkowski sum.
@@ IV-C2 Action
The action space $a$ is the goal of the trajectory which is composed of $(T_{target},d_{target},\sigma_{target},\dot{\sigma}_{target})$, where $T_{target}$ is the time interval between the current state and the goal state, and $\sigma_{target},d_{target}$ are the longitudinal and lateral target positions in the Frenet frame, and $\dot{\sigma}_{target}$ is the target longitudinal speed. Here, the trajectory planning method [28] is employed to plan a trajectory toward the goal. The lateral jerk-optimal trajectory is generated given the initial state of the AV $[d,\dot{d},\ddot{d}]$, and the end state $[d_{target},\dot{d}_{target}=0,\ddot{d}_{target}=0]$ at $T_{target}$ from the action. In addition, the longitudinal jerk-optimal trajectory is generated given the initial state of the AV $[\sigma,\dot{\sigma},\ddot{\sigma}]$, and the end state $[\sigma_{target},\dot{\sigma}_{target},\ddot{\sigma}_{target}=0]$ at $T_{target}$ from the action. The final trajectory is obtained by combining the lateral and longitudinal trajectories. Note that the planned trajectory is tracked using an MPC-based controller.
@@ IV-C3 Reward
where $\lambda_{lat\textunderscore acc}$, and $a_{lat\textunderscore acc}$ represent the weight and penalty for the lateral acceleration, $\lambda_{lat\textunderscore jerk}$, and $a_{lat\textunderscore jerk}$ are the weight and penalty for the lateral jerk, $\lambda_{long\textunderscore acc}$, and $a_{long\textunderscore acc}$ are the weight and penalty for the longitudinal acceleration, $\lambda_{long\textunderscore jerk}$, and $a_{long\textunderscore jerk}$ are the weight and penalty for the longitudinal jerk, $\lambda_{d}$, and $|d|$ are the weight and penalty for the lateral deviation from the target lane, $\lambda_{v}$, and $|v-v_{des}|$ are are the weight and the penalty for being slower or faster than the desired speed respectively, and $r_{collision}$ represents the reward and penalty for a collision event. Here, $r_{collision}$ is negative when a collision occurs and positive when no collision occurs. Note that the above reward is also utilized during RP.
@@ V Results
Figure 6: Average reward per time step in an episode, collision rate, success rate of all methods in every scenario during training.
Fig. 6, shows the average reward per time step in an episode, the collision rate, and the success rate of the compared methods for all scenarios during training. Table. I, shows the best scores of each of the compared methods for all scenarios during training. In Fig. 6, baseline 2 demonstrates greater learning stability than baseline 1 in every scenario. In terms of performance, in scenario 1, baseline 2 exhibits a $14.49\%$ lower collision rate and a $50.30\%$ higher success rate compared to baseline 1, in scenario 3, baseline 2 achieves an average reward of $0.1424$ that is $19.5$ times higher than that of baseline 1. Baseline 2 is more robust against action variance compared to baseline1, which enables baseline 2 to explore and identify optimal states and actions, leading to stable learning and improved performance. In Fig. 6, the proposed methods(i.e., the agent with RP, IRP, and IRP with uncertainty propagation) demonstrates stable learning compared to the baseline methods. In terms of performance, the proposed methods outperformed the baseline methods in all scenarios, with their performance improving sequentially. In scenario 2, the average reward of the RL agent with RP was $0.2200$, which was $5.7$ times higher than the baseline 1 and $2.4$ times higher than the baseline 2, the collision rate was $16.75\%$, which was $44.96\%$ lower than the baseline 1 and $26.29\%$ lower than the baseline2, and the success rate was $82.79\%$, which was $61.11\%$ higher than the baseline 1 and $26.06\%$ higher than the baseline 2. This increase in performance is attributed to the reduction of function approximation error by the RP method. Additionally, in scenario 4, the average reward of the RL agent with IRP was $0.2173$, which was $11.62\%$ higher than the agent with RP. The collision rate was $21.55\%$, which was $10.32\%$ lower than the agent with RP, and the success rate was $78.40\%$, which was $10.73\%$ higher than the agent with RP. The RL agent with IRP outperformed the agent with RP because IRP method predicts the agent’s action more accurately, leading to more precise reward predictions and better overall performance. The RL agent with IRP and uncertainty propagation performed even better than the agent with IRP due to the consideration of uncertainty. In scenario 1, the average reward of the RL agent with IRP and uncertainty propagation was $0.2297$, which was $4.69\%$ higher than the agent with IRP. The collision rate was $5.682\%$, which was $7.788\%$ lower than the agent with IRP, and the success rate was $93.86\%$, which was $9.05\%$ higher than the agent with IRP.
In scenario 3, even though lane changes are quite challenging in typical road situations, the collision rate was relatively small compared to the other scenarios. The reason for this is that the other traffic participants used in the experiment were designed to drive conservatively. The conservative driving of the other traffic participants reduced the collision rate. For scenario 4, the collision rate was much higher than in scenarios 1-3 because avoiding parked cars among the other traffic participants was considerably more challenging compared to the other scenarios. A video of the agent with IRP and uncertainty propagation method is available at https://www.youtube.com/watch?v=PfDbaeLfcN4.
TABLE I: Best results among the compared methods
@@ VI Conclusions And Future Work
In this study, we have proposed a method that includes RP, IRP, and uncertainty propagation methods to reduce the function approximation error of an AV’s RL-based planning agent. The proposed method was evaluated under several scenarios, and the results demonstrated that the proposed method improves both learning stability and agent performance compared to baseline methods. However, result of the evaluation showed that the proposed method still has poor performance in difficult and complex scenarios. Future works will involve increasing the safety while having better performance.

### v4

@@ III-D Uncertainty Propagation
During RP and IRP, the expected future states are considered; however, the variance and uncertainty of the predicted future states are not considered. Note that not considering the variance of the prediction leads to inaccurate function approximation and the RL agent being unaware of the uncertainty. In addition, ignoring the uncertainty can cause the agent to perform sudden deceleration or become involved in an accident. For example, if a low-functioning controller is employed to track a trajectory, and the RL agent does not take that into account, then the AV may encounter various dangerous situations and the predicted rewards may have significant error. To consider the variance of the prediction and the uncertainty of future states, uncertainty propagation on the RP and IRP motivated by [21, 22] is utilized.
The uncertainty propagation process is built on the Kalman filter; however, the measurement update process of the Kalman filter is removed because observing future states is impossible. The uncertainty propagation utilized on the RP is modeled as follows:
@@
where each state of the AV $\check{s}^{e}_{t,\tau}$, and states of the other traffic participants $\check{s}^{o}_{t,\tau}$ is modeled as Gaussian a random variable with means $s^{e}_{t,\tau}$, and $s^{o}_{t,\tau}$, and covariance of $\Sigma^{e}_{t,\tau}$, and $\Sigma^{o}_{t,\tau}$, respectively, based on the assumption made on III-A, $F$ is the state transition matrix and $Q^{e}_{\tau}$, $Q^{o}_{\tau}$ represent the process noise of the AV and the other traffic participants, respectively, where $Q^{e}_{\tau}$ is attributed to localization, control error, and $Q^{o}_{\tau}$ is attributed to object detection, trajectory prediction error respectively. The introduced states $\check{s}^{e}_{t,\tau}$ and $\check{s}^{o}_{t,\tau}$ are utilized in the collision checking process of the IRP method to account for uncertainty. Here, the ellipse defined by the covariance matrix can provide an upper bound of the probability $1-\delta$ that the AV and other traffic participants exist. However, to check for a collision between the AV and other traffic participants, the rectangle shapes of the AV and other traffic participants must be considered. In the proposed method, the Minkowski sum of the rectangle and the ellipse [21] is computed. The new shape from the Minkowski sum is then utilized for collision checking, which guarantees the probability of $(1-\delta)^{2}$ whether collide or not. The illustration of uncertainty propagation using the Minkowski sum is shown in Fig.1.
@@ III-E Overall Algorithm
Below, a demonstration showcasing the combination of IRP and uncertainty propagation is provided. The proposed method, based on the existing deterministic policy gradient algorithm, is applied during the critic update process. The pseudocode of the proposed method is given in Algorithm 1. Here, in lines 6-7, the uncertainty propagation and Minkowski sum are executed using the state of the AV and the other traffic participants $s^{e}_{t+\tau,0},s^{o}_{t+\tau,0}$. In line 9, the state $s_{t+\tau,\delta}$ results from the utilization of the agent’s previous state $s_{t+\tau,0}$, and the agent’s action $\pi(s_{t+\tau,0})$. In line 10, the reward is predicted during the transition from $s_{t+\tau,0}$ to $s_{t+\tau,\delta}$. In line 12, the state $s_{t+\tau,\delta}$ is divided into $s^{e}_{t+\tau,\delta}$, and $s^{o}_{t+\tau,\delta}$, which are used as the states $s^{e}_{t+\tau+\delta,0}$, and $s^{o}_{t+\tau+\delta,0}$ for the next iteration.
@@
3:     $r\leftarrow R_{t,\delta}$triangleright Initialize predicted return with received reward from the simulator
@@
10:         $R\leftarrow g(s,s^{\prime})$triangleright Predict the reward $R_{t+\tau,\delta}$
11:         $r\leftarrow r+\gamma^{\tau/\delta}R$triangleright Update the predicted return
12:         $s^{e},s^{o}\leftarrow f^{-1}(s^{\prime})$triangleright Update the next state
@@
17:     $s\leftarrow f({s^{e}}^{\prime},{s^{o}}^{\prime})$triangleright Merge states
18:     Set $y_{t}=r+\gamma^{T/\delta}Q(s,\pi(s|\theta^{\pi})|\theta^{Q})$
@@ IV Experiments
The proposed method and baseline methods were evaluated using the CARLA simulator. The experimental configuration, baseline methods, and implementation details are described in the following sections.
@@ IV-A Experiment Configuration
The proposed and baseline methods were evaluated in four distinct scenarios. Scenario 1 involved lane following with static obstacles, scenario 2 involved lane following with traffic participants, scenario 3 involved lane changing with traffic participants, and scenario 4 entailed overtaking parked cars with traffic participants. All necessary inputs, e.g., route path, object detection, trajectory prediction, and localization were given. Success was determined if the AV reached the goal without a collision within specified time. In this evaluation, the goal was $130m$ ahead of the AV’s initial position. Here, a maximum lateral deviation of $1.5m$ from the center of the target lane was permitted.
In each scenario, the AV was spawned on the road with a random lateral deviation of $[-1.5m,1.5m]$ from the center of the road, and a random heading angle deviation of $[-20deg,20deg]$ with a random initial speed of $[5km/h,15km/h]$. In scenario 1, a maximum of two static obstacles(i.e., vehicles) were spawned at random positions with a lateral deviation of $[-0.5m,0.5m]$ from the center of the road, and a heading angle deviation of $[-20deg,20deg]$. Scenario 2 included a maximum of five randomly spawned traffic participants. In scenario 3, the other traffic participants were the same as in scenario 2; however, the AV’s goal was to change lanes. In scenario 4, a maximum of two static obstacles(parked cars) and three traffic participants were spawned, and the goal of the AV was to overtake the parked cars while avoiding collisions with the other traffic participants. Note that the traffic participants in each scenario were designed to change lanes randomly.
@@ IV-B Baseline Methods
The following two baseline methods were considered in this evaluation. In baseline 1, the output of an agent was a control command, and it is identical to the method proposed in [5], except that in [5], monocular images are used as input. In baseline 2, the output of the agent was the trajectory goal, the same as that of the proposed method, for fair comparison. However, baseline 2 did not include RP, IRP, or uncertainty propagation. In addition, proposed method was evaluated individually as follows: RP, IRP, and IRP with uncertainty propagation.
@@ IV-C Implementation Details
All five methods, i.e., baseline1, baseline2, RP, IRP, and IRP with uncertainty propagation were implemented using the DDPG algorithm with the same state, and reward function.
@@ IV-C1 State
The features of the state space $s$ are composed of $s^{e}$ and $s^{o}$. Here, $s^{e}$ is the state of the AV and comprises $(d,\dot{d},\ddot{d},\dot{\sigma},\ddot{\sigma},\theta,v_{speed\_limit},l,w)$, where $\sigma$ and $d$ are the longitudinal and lateral position on the Frenet frame, $\theta$ is the heading angle difference with the center of the road, $v_{speed\_limit}$ is the speed limit of the road, and $l$ and $w$ are length and width of the AV in consideration of the Minkowski sum. $s^{o}$ is composed of $s^{o,k},k\in\mathbb{N}$, where $\mathbb{N}$ is a natural number. $s^{o,k}$ is the state of the $k$th traffic participant, which is composed of $(\sigma,d,\theta,l,w,v_{\sigma},v_{d})$, where $v_{\sigma}$ and $v_{d}$ are the longitudinal and lateral velocity on the Frenet frame.
@@ IV-C2 Action
The action space $a$ is the goal of the trajectory which is composed of $(T_{target},d_{target},\sigma_{target},\dot{\sigma}_{target})$, where $T_{target}$ is the time interval between the current state and the goal state, $\sigma_{target}$, and $d_{target}$ are the longitudinal and lateral target positions on the Frenet frame, and $\dot{\sigma}_{target}$ is the target longitudinal speed. Here, the trajectory planning method [28] is employed to plan a trajectory toward the goal. The lateral jerk-optimal trajectory is generated given the initial state of the AV $[d,\dot{d},\ddot{d}]$, and the end state $[d_{target},\dot{d}_{target}=0,\ddot{d}_{target}=0]$ at $T_{target}$ from the action. In addition, the longitudinal jerk-optimal trajectory is generated given the initial state of the AV $[\sigma,\dot{\sigma},\ddot{\sigma}]$, and the end state $[\sigma_{target},\dot{\sigma}_{target},\ddot{\sigma}_{target}=0]$ at $T_{target}$ from the action. The final trajectory is obtained by combining the lateral and longitudinal trajectories. Note that the planned trajectory is tracked using an MPC-based controller.
@@ IV-C3 Reward
where $\lambda_{lat\textunderscore acc}$, and $a_{lat\textunderscore acc}$ represent the weight and penalty for the lateral acceleration, $\lambda_{lat\textunderscore jerk}$, and $a_{lat\textunderscore jerk}$ are the weight and penalty for the lateral jerk, $\lambda_{long\textunderscore acc}$, and $a_{long\textunderscore acc}$ are the weight and penalty for the longitudinal acceleration, $\lambda_{long\textunderscore jerk}$, and $a_{long\textunderscore jerk}$ are the weight and penalty for the longitudinal jerk, $\lambda_{d}$, and $|d|$ are the weight and penalty for the lateral deviation from the target lane, $\lambda_{v}$, $|v-v_{des}|$ are the weight and the penalty for being slower or faster than the desired speed respectively, and $r_{collision}$ represents the reward and penalty for a collision event. Here, $r_{collision}$ is negative when a collision occurs and positive when no collision occurs. Note that the above reward is also utilized during RP.
@@ V Results
Figure 5: Average reward per time step in an episode, collision rate, and success rate of all methods for all scenarios during training.
The average reward per time step, collision rate, and success rate across all training scenarios are depicted in Fig. 5. Table I showcases the best scores achieved by each method across all training scenarios, providing a comprehensive overview of performance comparisons.
Baseline 2 consistently outperformed baseline 1 in every scenario and key metric while demonstrating stable learning. The robustness of baseline 2 against action variance enabled it to explore and identify optimal states and actions, resulting in stable learning and improved performance. Furthermore, agents utilizing RP, IRP, and IRP with uncertainty propagation exhibit stable learning compared to the baseline methods, as illustrated in Fig. 5. Across all scenarios, these agents consistently outperformed the baseline methods, showing sequential performance improvements. The agent with RP outperformed the baseline methods, which is attributed to its utilization of the error reduction property of N-step SARSA and explicit learning of the consequences of actions. The agent with IRP surpassed the agent with RP due to accurate prediction of future actions, resulting in more precise reward prediction and improved overall performance. Additionally, the agent with IRP and uncertainty propagation outperformed the agent with IRP, benefiting from the consideration of uncertainty.
In scenario 3, although lane changes are quite challenging in typical road situations, the collision rate was relatively low compared to the other scenarios. This was due to the conservative driving behavior of the other traffic participants used in the experiment, which helped reduce the collision rate. In scenario 4, the collision rate was much higher than in scenarios 1-3 because it was considerably more challenging to avoid parked cars while also avoiding the other traffic participants, compared to the other scenarios. A video of the agent with IRP and uncertainty propagation is available at https://www.youtube.com/watch?v=PfDbaeLfcN4.
TABLE I: Best scores for each compared method for all scenarios during training.
@@ VI Conclusions And Future Work
In this study, methods including RP, IRP, and uncertainty propagation are proposed to reduce the function approximation error and improve the performance and learning stability of an AV’s RL-based planning agent. The proposed method was evaluated under several scenarios, and the results demonstrated that the proposed method improves both learning stability and agent performance compared to baseline methods. However, the result of the evaluation showed that the proposed method still has poor performance in difficult and complex scenarios. Future work will involve increasing safety while having better performance.